In what does the x-axis divides the line segement joining the points (2,-3) and (5,6)

Grade:11

1 Answers

Vikas TU
7 years ago
Eqn. of linee joining these two points is:
(y – 6) = ((-3-6)/(2-5))(x – 5)
y – 6 = (-9/-3)(x – 5)
y – 6 = 3x – 15
y = 3x – 9
when y = 0
3x = 9 
x = 3
at x = 3 it cuts the x – axis.